iii) The sum of the digits of a two-digit number is 9. Also, nine times this number is twice the number obtained by reversing the order of the digits. Find the number.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Let the unit digit and tens digit of a number be x and y respectively. Then, Number (n) = 10B + A N after reversing order of the digits = 10A + B According to the given information, A + B = 9…………………….(i) 9(10B + A) = 2(10A + B) 88 B – 11 A = 0 -A + 8B = 0 ………………………………………………………….. (ii) Adding the equations (i) and (ii) we get, 9B = 9 B = 1……………………………………………………………………….(3) Substituting this value of B, in the equation (i) we get A= 8 Hence the number (N) is 10B + A = 10 x 1 +8 = 18